In the chemical reaction below, how many grams of KCN will react with 81.55g of H2SO4? Show the solution.




2KCN + H2SO4 ------> K2SO4 + 2HCN

1
Expert's answer
2022-06-03T06:24:04-0400

n(H2SO4)=81.55/98=0.832 mol, m(KCN)=65*0.832*2=108.2